# Embers of Active Galactic Nuclei: Tidal Disruption Events and Quasi-periodic Eruptions

### v1

@@ Abstract
Recent observations have confirmed the direct association between tidal disruption events (TDEs) and quasi-periodic eruptions (QPEs). In addition, TDE hosts and QPE hosts are statistically found to be similar in their morphological properties and in the strong overrepresentation of post-starburst galaxies. Particularly, both of them show an intriguing preference for extending emission line regions (EELRs), which are indicative of recently faded active galactic nuclei (AGNs). This further suggests that QPEs might be produced following TDEs involving supermassive black holes at a particular stage, when the AGN activity has recently ceased. Moreover, in the framework of “QPEs=EMRI+accretion disk” model, a large fraction of QPE EMRIs are inferred to be quasi-circular from the QPE timing, indicating that they are wet EMRIs that were formed in the AGN disk during a previous AGN phase. Based on these facts, we propose a unified scenario that connects these three phenomena: AGN activities boost both the TDE rate and the formation rate of low-eccentricity EMRIs, consequently TDEs are preferentially found in recently faded AGNs instead of in on-going AGNs due to selection effects, and QPEs are also preferentially found in recently faded AGNs where TDEs frequently feed a misaligned accretion disk to the EMRI.
@@ 1 Introduction
In the past decade, advancements in survey technologies across various wavelengths have resulted in a fast increasing detection rate of rare transient events associated with supermassive black holes (SMBHs) located in the centers of galaxies. They include stellar tidal disruption events (TDEs, see Gezari 2021 as a review), changing-look active galactic nuclei (AGNs, LaMassa et al. 2015) and X-ray quasi-periodic eruptions (QPEs, Miniutti et al. 2019). Among them the physical origin of QPEs is particularly puzzling due to their fascinating observational properties and rarity, with about 10 QPE sources having been reported so far (Sun et al., 2013; Miniutti et al., 2019; Giustini et al., 2020; Arcodia et al., 2021, 2022, 2024; Chakraborty et al., 2021; Evans et al., 2023; Guolo et al., 2024; Nicholl et al., 2024). A significant progress very recently is the direct detection of QPEs following a confirmed optical TDE AT2019qiz (Nicholl et al., 2024), indicating that at least some (if not all) QPEs are firmly connected with TDEs. Such a connection was previously tentatively observed in the first QPE source GSN 069 (Miniutti et al., 2019) although it is not a standard TDE with optical spectroscopic identification yet with evidence in its long-term X-ray light curve (Shu et al., 2018) and abnormal nitrogen-enriched gas implied by its UV and X-ray spectra (Sheng et al., 2021; Kosec et al., 2025). In addition, the QPE-TDE association also appears in some other sources prior to AT 2019qiz, although it is not as conclusive, due to a lack of clear confirmation of QPE, i.e. AT 2019vcb (Quintin et al., 2023), or TDE, i.e. XMMSL1 J024916.6-041244 (Chakraborty et al., 2021). Additional eruptions of AT 2019vcb reported by Bykov et al. (2024) further increase its credibility as a QPE source.
@@
Further insights can be gained from the resemblance in host galaxy characteristics between TDEs and QPEs. TDEs are found to be strongly overrepresented in low-mass post-starburst (PSB) or E+A galaxies (Arcavi et al., 2014; French et al., 2016; Law-Smith et al., 2017) which show a highly concentrated bulge with high Sérsic indices, bulge-to-total light (B/T) ratios, and stellar surface mass densities compared to the overall galaxy population (French et al., 2020). Interestingly, a similar preference of host galaxies has been also observed in QPEs recently (Wevers et al., 2022; Gilbert et al., 2024). In addition, Wevers et al. (2022) performed a systematic study of the host galaxies of five known QPE sources using new and archived medium-resolution optical spectra and claimed that all QPE host galaxies have emission lines indicating the presence of an AGN. Further integral field spectroscopic observations suggest a high incidence (3/5) of extended emission line region (EELR) detection in the hosts of QPEs (Wevers et al., 2024). These EELRs extend up to 10 kpc in size, requiring photoionisation by a nonstellar continuum, but far beyond the current AGN luminosity. In a companion study, Wevers & French (2024) estimated the incidence of EELRs in PSB-TDE hosts and found it is a factor of 10$\sim$ 10$\times$ 10 × higher than in other PSB galaxies, suggesting that a gas-rich postmerger environment is a key ingredient in driving elevated TDE rates. Therefore, both TDEs and QPEs actually show a preference for a specific evolutionary stage of PSB Galaxies, namely the phase of recently faded AGNs (French et al., 2023).
@@
The growing evidence of TDE-QPE association promoted a “TDE+EMRI=QPEs” model (Linial & Metzger, 2023), where the QPEs are the result of quasi-periodic collisions between an EMRI (extreme mass ratio inspiral) and an accretion disk formed from a previous TDE. This model or the general EMRI+accretion disk model (Franchini et al., 2023; Tagawa & Haiman, 2023) naturally explains the intriguing features in the QPE timing, including the alternating long-short $T_{\rm long}-T_{\rm short}$ pattern found in recurrence times of several QPE sources (Miniutti et al., 2019; Giustini et al., 2020; Arcodia et al., 2021, 2022; Miniutti et al., 2025), and an approximately constant sum of two consecutive recurrence times $T_{\rm long}+T_{\rm short}$ with both $T_{\rm long}$ and $T_{\rm short}$ showing large variations with time (Zhou et al., 2024a). In the framework of EMRI+accretion disk model, previous studies have demonstrated the feasibility of measuring the SMBH mass and the EMRI orbital parameters from the QPE timing (Xian et al., 2021; Franchini et al., 2023; Zhou et al., 2024a, c, b). Systematic analyses show that EMRIs that are sourcing QPEs are with orbital radii of $\mathcal{O}(10^{2})M_{\bullet}$, and a large fraction of them are of low orbital eccentricity with $e=\mathcal{O}(10^{-2})$ (Zhou et al., 2024a, c, b), where $M_{\bullet}$ is the gravitational radius of the SMBH. These low-eccentricity QPE EMRIs can be naturally interpreted as wet EMRIs formed in accretion disks of previous AGNs (Sigl et al., 2007; Levin, 2007; Pan & Yang, 2021a; Pan et al., 2021; Pan & Yang, 2021b; Pan et al., 2022; Derdzinski & Mayer, 2023), but are too circular to be compatible with predictions of loss-cone channel (Hopman & Alexander, 2005; Preto & Amaro-Seoane, 2010; Bar-Or & Alexander, 2016; Babak et al., 2017; Amaro-Seoane, 2018; Broggi et al., 2022) or Hills channel (Miller et al., 2005; Raveh & Perets, 2021). Therefore, past AGN activities are inevitable for providing the necessary condition (a low-eccentricity EMRI) for a large fraction of QPEs.
@@ 2 QPE Sample
Note. — Columns (1)-(3): the name, redshift and black hole mass in units of $\log_{10}(M_{\bullet}/M_{\odot})$ of the QPE sources. Columns (4)-(5): Whether or not the QPE is associated with a TDE or a recently faded AGN. Here ”?” means that there is no clear evidence yet, but this may be due to a lack of sufficient observations. Column (6): Related references: 1 (Wevers et al., 2022); 2 (Yao et al., 2023); 3 (Arcodia et al., 2024); 4 (Wevers et al., 2024); 5 (Shu et al., 2018); 6 (Sheng et al., 2017); 7 (Nicholl et al., 2024); 8 (Xiong et al. in prep.); 9 (Arcodia et al., 2021); 10 (Chakraborty et al., 2021); 11 (Quintin et al., 2023); 12 (Bykov et al., 2024). Note the black hole mass measurements in column (3) depend on the scaling relation used, e.g., a different $M_{\bullet}-\sigma_{\star}$ relation used in Wevers et al. (2019) prefers a lower value of J0249 black hole mass $\log_{10}(M_{\bullet}/M_{\odot})=4.93^{+0.55}_{-0.53}$.
@@
  • GSN 069: This is the first confirmed QPE source (Miniutti et al., 2019). It was initially reported as a true Seyfert 2 galaxy candidate since a bright soft X-ray source was detected at a position consistent with it in 2010, but its optical spectra are always dominated by unresolved emission lines with no broad components (Miniutti et al., 2013). Its nondetection by ROSAT implies it has undergone $\gtrsim 240$ times of X-ray brightening, suggesting a likely TDE, which has been supported by its subsequent long-term slow decay (Shu et al., 2018). The VLT/MUSE observation reveals a EELR with $R_{\rm max}\sim 9$ kpc, which calls for a $L_{\rm ion,min}$ of $6\times 10^{42}$ erg s-1 (Wevers et al., 2024). It is also notable that Patra et al. (2024) found evidence of a compact nuclear [O III] region making use of archival HST data, which indicates that the accretion activity is only $\sim 10-100$ yr old.
  • AT2019qiz: This is the first optical TDE with unambiguous QPE detection (Nicholl et al., 2024). The ratios of narrow lines in the BPT diagram lie intermediate between the region of star-forming galaxies and AGNs, indicative of a possible weak AGN prior to the TDE (Nicholl et al., 2020). After a careful study of the VLT/MUSE data of its host, we found an EELR, which is most prominent in the N II map with scale of $\sim 3.7$ kpc. It provides further evidence of AGN ionizing in the near past (Xiong et al. in prep.). The detection of a prominent IR echo (Short et al., 2023) also suggests a likely residual dusty torus of a recently faded AGN. In contrast, the echoes of normal TDEs are rather weak due to a much lower dust covering factor (Jiang et al., 2021).
@@
  • XMMSL1 J024916.6-041244J0249 (hereafter J0249): This is the fifth QPE candidate with 1.5 QPE-like soft X-ray flares discovered by a blind search of the XMM-Newton Source Catalog (Chakraborty et al., 2021). It was first identified as a TDE candidate based on a flux increase by a factor of 88 compared to a ROSAT upper limit, and a very soft X-ray spectrum that can be described by a blackbody (Strotjohann et al., 2016). Its host galaxy was initially classified as a Seyfert 1.9 (Esquej et al., 2007), which has recently been reclassified as a star-forming galaxy potentially hosting an AGN by higher resolution spectra (Wevers et al., 2019, 2022).
@@ 4 Conclusion: a unified scenario
As depicted in Figure 1, the AGN phase plays a central role in this scenario. In the standard AGN phase (panel a), both the TDE rate and the EMRI formation rate are boosted. However, observationally, the identification of TDEs in AGNs is rather complicated due to the dilution of bright AGN emission and the lack of predicted hallmark features for these special TDEs. As the AGN phase is off, the accretion disk has vanished and an EMRI orbiting around the SMBH at radius $r=\mathcal{O}(10^{2})M_{\bullet}$ is left (panel b). A TDE then occurs during the recently faded AGN phase, when the TDE rate remains elevated and there is no challenge to identify it without AGN (panel c). After $\mathcal{O}(1)$ years, a TDE disk forms which is generally misaligned with respect to the EMRI, as a result, QPEs are generated from quasi-periodic collisions between the EMRI and the disk (panel d).
@@
We thank the referee for his/her very positive and constructive comments, which have improved the manuscript significantly. This work is supported by the National SKA Program of China (2022SKA0130102), the National Natural Science Foundation of China (grants 12192221,12393814), the Strategic Priority Research Program of the Chinese Academy of Sciences (XDB0550200) and the China Manned Space Project.

### v2

# Embers of Active Galactic Nuclei: Tidal Disruption Events and Quasiperiodic Eruptions

@@ Abstract
Recent observations have confirmed the direct association between tidal disruption events (TDEs) and quasiperiodic eruptions (QPEs). In addition, TDE hosts and QPE hosts are statistically found to be similar in their morphological properties and in the strong overrepresentation of poststarburst galaxies. Particularly, both of them show an intriguing preference for extending emission line regions (EELRs), which are indicative of recently faded active galactic nuclei (AGNs). This further suggests that QPEs might be produced following TDEs involving supermassive black holes at a particular stage, when the AGN activity has recently ceased. Moreover, in the framework of “QPEs=extreme mass ratio inspiral (EMRI) + accretion disk” model, a large fraction of QPE EMRIs are inferred to be quasi-circular from the QPE timing, indicating that they are wet EMRIs that were formed in the AGN disk during a previous AGN phase. Based on these facts, we propose a unified scenario that connects these three phenomena: AGN activities boost both the TDE rate and the formation rate of low-eccentricity EMRIs, consequently TDEs are preferentially found in recently faded AGNs instead of in ongoing AGNs due to selection effects, and QPEs are also preferentially found in recently faded AGNs where TDEs frequently feed a misaligned accretion disk to the EMRI.
@@ 1 Introduction
In the past decade, advancements in survey technologies across various wavelengths have resulted in a fast increasing detection rate of rare transient events associated with supermassive black holes (SMBHs) located in the centers of galaxies. They include stellar tidal disruption events (TDEs, see Gezari 2021 as a review), changing-look active galactic nuclei (AGNs, LaMassa et al. 2015) and X-ray quasiperiodic eruptions (QPEs, Miniutti et al. 2019). Among them the physical origin of QPEs is particularly puzzling due to their fascinating observational properties and rarity, with about 10 QPE sources having been reported so far (Sun et al., 2013; Miniutti et al., 2019; Giustini et al., 2020; Arcodia et al., 2021, 2022, 2024; Chakraborty et al., 2021; Evans et al., 2023; Guolo et al., 2024; Nicholl et al., 2024). A significant progress very recently is the direct detection of QPEs following a confirmed optical TDE AT2019qiz (Nicholl et al., 2024), indicating that at least some (if not all) QPEs are firmly connected with TDEs. Such a connection was previously tentatively observed in the first QPE source GSN 069 (Miniutti et al., 2019) although it is not a standard TDE with optical spectroscopic identification yet with evidence in its long-term X-ray light curve (Shu et al., 2018) and abnormal nitrogen-enriched gas implied by its UV and X-ray spectra (Sheng et al., 2021; Kosec et al., 2025). In addition, the QPE-TDE association also appears in some other sources prior to AT 2019qiz, although it is not as conclusive due to a lack of clear confirmation of QPE, in the case of AT 2019vcb (Quintin et al., 2023), or TDE, in the case of XMMSL1 J024916.6-041244 (Chakraborty et al., 2021). Additional eruptions of AT 2019vcb reported by Bykov et al. (2024) further increase its credibility as a QPE source.
@@
Further insights can be gained from the resemblance in host galaxy characteristics between TDEs and QPEs. TDEs are found to be strongly overrepresented in low-mass poststarburst (PSB) or E+A galaxies (Arcavi et al., 2014; French et al., 2016; Law-Smith et al., 2017) which show a highly concentrated bulge with high Sérsic indices, bulge-to-total light (B/T) ratios, and stellar surface mass densities compared to the overall galaxy population (French et al., 2020). Interestingly, a similar preference of host galaxies has been also observed in QPEs recently (Wevers et al., 2022; Gilbert et al., 2024). In addition, Wevers et al. (2022) performed a systematic study of the host galaxies of five known QPE sources using new and archived medium-resolution optical spectra and claimed that all QPE host galaxies have emission lines indicating the presence of an AGN. Further integral field spectroscopic observations suggest a high incidence (3/5) of extended emission line region (EELR) detection in the hosts of QPEs (Wevers et al., 2024). These EELRs extend up to 10 kpc in size, requiring photoionization by a nonstellar continuum, but far beyond the current AGN luminosity. In a companion study, Wevers & French (2024) estimated the incidence of EELRs in PSB-TDE hosts and found it is a factor of 10$\sim$ 10$\times$ 10 × higher than in other PSB galaxies, suggesting that a gas-rich postmerger environment is a key ingredient in driving elevated TDE rates. Therefore, both TDEs and QPEs actually show a preference for a specific evolutionary stage of PSB galaxies, namely the phase of recently faded AGNs (French et al., 2023).
@@
The growing evidence of TDE-QPE association promoted a “TDE+EMRI=QPEs” model (Linial & Metzger, 2023), where the QPEs are the result of quasiperiodic collisions between an extreme mass ratio inspiral (EMRI) and an accretion disk formed from a previous TDE. This model or the general EMRI+accretion disk model (Franchini et al., 2023; Tagawa & Haiman, 2023) naturally explains the intriguing features in the QPE timing, including the alternating long-short $T_{\rm long}-T_{\rm short}$ pattern found in recurrence times of several QPE sources (Miniutti et al., 2019; Giustini et al., 2020; Arcodia et al., 2021, 2022; Miniutti et al., 2025), and an approximately constant sum of two consecutive recurrence times $T_{\rm long}+T_{\rm short}$ with both $T_{\rm long}$ and $T_{\rm short}$ showing large variations with time (Zhou et al., 2024a). In the framework of EMRI+accretion disk model, previous studies have demonstrated the feasibility of measuring the SMBH mass and the EMRI orbital parameters from the QPE timing (Xian et al., 2021; Franchini et al., 2023; Zhou et al., 2024a, c, b). Systematic analyses show that EMRIs that are sourcing QPEs are with orbital radii of $\mathcal{O}(10^{2})M_{\bullet}$, and a large fraction of them are of low orbital eccentricity with $e=\mathcal{O}(10^{-2})$ (Zhou et al., 2024a, c, b), where $M_{\bullet}$ is the gravitational radius of the SMBH. These low-eccentricity QPE EMRIs can be naturally interpreted as wet EMRIs formed in accretion disks of previous AGNs (Sigl et al., 2007; Levin, 2007; Pan & Yang, 2021a; Pan et al., 2021; Pan & Yang, 2021b; Pan et al., 2022; Derdzinski & Mayer, 2023), but are too circular to be compatible with predictions of loss-cone channel (Hopman & Alexander, 2005; Preto & Amaro-Seoane, 2010; Bar-Or & Alexander, 2016; Babak et al., 2017; Amaro-Seoane, 2018; Broggi et al., 2022) or Hills channel (Miller et al., 2005; Raveh & Perets, 2021). Therefore, past AGN activities are inevitable for providing the necessary condition (a low-eccentricity EMRI) for a large fraction of QPEs.
@@ 2 QPE Sample
Note. — Columns (1)-(3): the name, redshift and black hole mass in units of $\log_{10}(M_{\bullet}/M_{\odot})$ of the QPE sources. Columns (4)-(5): Whether or not the QPE is associated with a TDE or a recently faded AGN. Here ”?” means that there is no clear evidence yet, but this may be due to a lack of sufficient observations. Column (6): Related references: 1 (Wevers et al., 2022); 2 (Yao et al., 2023); 3 (Arcodia et al., 2024); 4 (Wevers et al., 2024); 5 (Shu et al., 2018); 6 (Sheng et al., 2017); 7 (Nicholl et al., 2024); 8 (Xiong et al., 2025); 9 (Arcodia et al., 2021); 10 (Chakraborty et al., 2021); 11 (Quintin et al., 2023); 12 (Bykov et al., 2024). Note the black hole mass measurements in column (3) depend on the scaling relation used, e.g., a different $M_{\bullet}-\sigma_{\star}$ relation used in Wevers et al. (2019) prefers a lower value of J0249 black hole mass $\log_{10}(M_{\bullet}/M_{\odot})=4.93^{+0.55}_{-0.53}$. Also we noticed that the mass of eRO-QPE4 might be overestimated systematically.
@@
  • GSN 069: This is the first confirmed QPE source (Miniutti et al., 2019). It was initially reported as a true Seyfert 2 galaxy candidate since a bright soft X-ray source was detected at a position consistent with it in 2010, but its optical spectra are always dominated by unresolved emission lines with no broad components (Miniutti et al., 2013). Its nondetection by ROSAT implies it has undergone X-ray brightening by a factor of $\gtrsim 240$, suggesting a likely TDE, which has been supported by its subsequent long-term slow decay (Shu et al., 2018). The VLT/MUSE observation reveals a EELR with $R_{\rm max}\sim 9$ kpc, which calls for a $L_{\rm ion,min}$ of $6\times 10^{42}$ erg s-1 (Wevers et al., 2024). It is also notable that Patra et al. (2024) found evidence of a compact nuclear [O III] region making use of archival HST data, which indicates that the accretion activity is only $\sim 10-100$ yr old.
  • AT2019qiz: This is the first optical TDE with unambiguous QPE detection (Nicholl et al., 2024). The ratios of narrow lines in the BPT diagram lie intermediate between the region of star-forming galaxies and AGNs, indicative of a possible weak AGN prior to the TDE (Nicholl et al., 2020). After a careful study of the VLT/MUSE data of its host, we found an EELR, which is most prominent in the N II map with scale of $\sim 3.7$ kpc. It provides further evidence of AGN ionizing in the near past (Xiong et al., 2025). The detection of a prominent IR echo (Short et al., 2023) also suggests a likely residual dusty torus of a recently faded AGN. In contrast, the echoes of normal TDEs are rather weak due to a much lower dust covering factor (Jiang et al., 2021).
@@
  • XMMSL1 J024916.6-041244 (hereafter J0249): This is the fifth QPE candidate with 1.5 QPE-like soft X-ray flares discovered by a blind search of the XMM-Newton Source Catalog (Chakraborty et al., 2021). It was first identified as a TDE candidate based on a flux increase by a factor of 88 compared to a ROSAT upper limit, and a very soft X-ray spectrum that can be described by a blackbody (Strotjohann et al., 2016). Its host galaxy was initially classified as a Seyfert 1.9 (Esquej et al., 2007), which has recently been reclassified as a star-forming galaxy potentially hosting an AGN by higher resolution spectra (Wevers et al., 2019, 2022).
@@ 4 Conclusion: a unified scenario
As depicted in Figure 1, the AGN phase plays a central role in this scenario. In the standard AGN phase (panel a), both the TDE rate and the EMRI formation rate are boosted. However, observationally, the identification of TDEs in AGNs is rather complicated due to the dilution of bright AGN emission and the lack of predicted hallmark features for these special TDEs. As the AGN phase is off, the accretion disk has vanished and an EMRI orbiting around the SMBH at radius $r=\mathcal{O}(10^{2})M_{\bullet}$ is left (panel b). A TDE then occurs during the recently faded AGN phase, when the TDE rate remains elevated and there is no challenge to identify it without AGN (panel c). After $\mathcal{O}(1)$ years, a TDE disk forms which is generally misaligned with respect to the EMRI, as a result, QPEs are generated from quasiperiodic collisions between the EMRI and the disk (panel d).
@@
We thank the referee for very positive and constructive comments, which have improved the manuscript significantly. This work is supported by the National SKA Program of China (2022SKA0130102), the National Natural Science Foundation of China (grants 12192221,12393814), the Strategic Priority Research Program of the Chinese Academy of Sciences (XDB0550200) and the China Manned Space Project.